# Plots unlock time-series understanding in multimodal models

### v1

@@ 1 Introduction
Multimodal models like GPT4 (Achiam et al., 2023) and Gemini (Gemini Team et al., 2023) are trained to understand visual information natively. However, they are not specifically trained to understand time-series data – in particular, the tokenizers for large language models (LLMs) are not well-suited for representing large sequences of floating point numbers (Spathis & Kawsar, 2024). This mirrors the human approach; we cannot easily make sense of a long array of floating point numbers - instead our first instinct is often to visualize the data through plotting, followed by extracting insights through statistical analysis.
@@ 2 Related Work
We use the term “time-series understanding” in this paper to distinguish from time-series forecasting. Time-series forecasting predicts future data points based on points seen so far, whereas we are primarily interested in the setting where we connect the time-series data to a multimodal model for further analysis. In particular, we want to show that multimodal models can reason about overall trends, the relationship between multiple time-series, overall clustering of data, and other time-series understanding tasks. For a closer look at the literature, the survey paper by Zhang et al. (2024) tracks a wide range of time-series understanding and forecasting work.
We deliberately avoid using multimodal models to generate any of the questions themselves (e.g. SPIQA (Pramanick et al., 2024), CharXiv (Wang et al., 2024)) as this can introduce biases during evaluation that are hard to account for (e.g. favoring their own output as in (Panickssery et al., 2024)).
LLMTime (Gruver et al., 2024) shows that with careful tokenization, text-only LLMs can perform well at forecasting tasks. We perform ablations based on their methods and select the best tokenizations accordingly for our text baselines.
@@
There are several existing approaches that train time-series encoders for specific tasks or domains. For example, Chan et al. (2024) train a domain-specific encoder for multimodal medical time-series analysis. Similarly, Cosentino et al. (2024) train an encoder for the sleep data in their Digital Well-being task. In this paper, we are not claiming that our approach would outperform a task-specific encoder on a specific task – we claim that one can achieve much better performance from a foundation model by exploiting its native multimodal capabilities compared to using only text. We want to show that simply plotting the data is at the very least an easy first step, and might be a helpful approach when training a task-specific encoder from scratch may not be feasible due to the requirements on having additional paired data, compute and expertise.
Closely related to our work are those methods that use vision-embedding models like Contrastive Language–Image Pre-training (CLIP) by Radford et al. (2021). Wimmer & Rekabsaz (2023) use CLIP to embed plots of financial time-series data, from which features are extracted for use by downstream classifiers. Since we use the vision encoders of the multimodal foundation models directly, there is no need for further feature extraction or downstream classifiers in our approach. IMU2CLIP (Moon et al., 2023) and ImageBind (Girdhar et al., 2023) use video and image data paired with waveforms to learn joint embeddings. Both these works rely on existing paired waveform and video data to “bind” the modalities together, whereas we can simply plot the waveforms and use the existing multimodal vision encoder to derive our time-series embeddings.
Perhaps an inversion of our approach, DePlot (Liu et al., 2023) translates plots to tables and operates on the tabular data. This approach may be sound for discrete data where the number of points remains small – cases where a human would be expected to understand a table of data well.
@@
Finally, Rahmanzadehgervi et al. (2024) suggest that multimodal models are unable to reason about visual inputs, although some follow-up work by Corin (2024) indicates that prompt engineering can fix losses. In any case, our results do not necessarily contradict this - for many of our tasks humans may be able to get perfect scores, and indeed the multimodal models do not. Regardless, our main claim that plots are better suited than text as input to a multimodal model for time-series understanding holds true.
@@ 3 Methodology
We evaluate our visual prompting method on both synthetic data and real-world use-cases. Synthetic data allows us to control the difficulty of the task by adding noise and altering the number of data points, and to investigate specific kinds of reasoning in isolation. We chose the synthetic tasks to align with the different steps of reasoning we hypothesize are required for the representative real-world use cases we test on. These include understanding the local and global longitudinal signatures (trend and magnitude) of a time-series, and potentially comparing it with several other time-series (as in the case of multidimensional sensing). We summarize in Section 4.1 the different tasks in our experiments, along with the type of reasoning we are probing.
@@ 4 Experiments and Results
In this section we provide a summary of each task and the differences in plot- versus text-based model performances. Table 1 summarizes the tasks and the reasoning required for each, and Table 2 summarizes the model performances. More details about each task can be found in Supplementary Section S.1.1.
Type of data Task Reasoning Num points
@@
Table 1: Summary of the eight tasks we study in this paper, including the type of reasoning each requires and the scale of the number of points in the time-series.
@@
Shots GPT4o-mini Gemini Flash 1.5 GPT4o Gemini Pro 1.5
@@
Table 2: Summary of results across all tasks. Rows are different tasks, potentially with different numbers of few-shot examples, and columns are the gains in using plots over texts for the specified models for the specified metric (a positive number always means plots are better). Cell entries contain the median and inter-quartile range (IQR) of the metrics as described in Section 3.4. Bold values are a visual clue that the plot method performs better than the text as judged by a positive median difference between the two performances. Stars in the synthetic tasks indicate statistically significant differences between the plot and text metrics at 95% confidence corrected for multiple comparisons; we could not perform the same hypothesis testing on the real-world tasks due to non-independence of bootstrapped samples (see Section 3.4 for more information). See Supplementary Table S1 for relative differences between plot and task approaches.
@@ 4.1 Synthetic Data Tasks
Figure 1 summarizes all the zero-shot versions of the synthetic data tasks showing that plot-based methods outperform the text-based methods across GPT and Gemini model families, with few exceptions.
@@
Figure 1: Zero-shot synthetic data results showing plot- and text-based accuracy (or MAE for the cluster counting task) distributions for all models. Also shown is a horizontal line representing random performance.
@@
This is the simplest task that requires only identifying one overall trend and correctly classifying it into one of five functional tasks (linear, quadratic, cubic, exponential or periodic).
@@
This task now requires understanding the trends in two lines and comparing them against each other to correct identify whether the lines are positively or negatively correlated.
@@
In this task, the model needs to correctly identify the $N$ number of clusters present in a set of points.
@@
This is a harder task: the model must now identify the correct first derivative (out of four choices) of the function provided in the question. The function and choices are presented as either plots or text. We pass various known functions to the model alongside four synthetic first derivatives, each corresponding to different functional classes, and ask it to identify which of the multiple choices corresponds to the true derivative.
@@
As a hard variant of derivative identification, the model must now identify the correct linear function (out of four choices) that corresponds to the quadratic function in the question. Here the model must pay attention to both the sign and magnitude of the slope.
@@
Figure 2: Few-shot quadratic derivative identification
@@ 4.2 Real-world Tasks
Figure 3: Results of fall detection task for all models, for 1, 3, 5 and 10 few-shots. We also report that the best performing plot models on 10-shot have (sensitivity, specificity) as follows: Gemini Pro 1.5 - (0.84, 0.95) and GPT4o - (0.92, 0.81). The state-of-the-art result reported by Aziz et al. (2017) on this dataset is a purpose-built support vector machine (SVM) trained on 50% of the data which achieves (0.96, 0.96).
An IMU segment is a 6D-vector composed of 3-axes accelerometer signals and 3-axes gyroscope signals. The first real-world task we evaluate is to classify whether a 15-second IMU segment recorded at 128hz contains a fall, a “near” fall or showed “active daily living” (ADL). The dataset used in the open-source IMU Fall Detection Dataset (IMUFD, Aziz et al. (2017)).
@@
Figure 4: Results of activity detection task for all models, for 1, 3, 5 and 10 (where context length allowed) few-shots. The state-of-the-art result reported by Kumar & Selvam (2022) is an average $F_{1}$ score evaluating their deep learning activity recognition model on HHAR.
A further real-world IMU task we evaluate is to classify whether a 15-second IMU segment is one of five activity classes: “sit”, “stand”, “stairs”, “walk” or “bike”. The dataset used in the open-source Heterogeneity Human Activity Recognition dataset (HHAR, Stisen et al. (2015)). As with Fall Detection we test performance with 1, 3, 5 and 10 few-shot examples, with 383 samples per model and number of few-shots. For this task the text representation of the 10 few-shot examples exceeded the GPT4o and GPT4o-mini 128k context windows, so these results are only shown for the Gemini models.
@@
Figure 5: Results of readiness task for Gemini models only (as the dataset cannot be sent to other models).
@@
We use training load data from 350 fitness case studies (Cosentino et al., 2024) and present it as tables or TRIMP bar plots. Each case study contains data from 30 consecutive days. We use a simplified version of the textual prompt and visualization from Cosentino et al. (2024) and do not split TRIMP in different heart rate zones. We tested Gemini 1.5 Pro and Gemini 1.5 Flash for both plot and text approaches as zero-shot tasks. Since this task involved analyzing just 30 data points we did not expect the plot prompt to excel here. Interestingly, models of different sizes showed opposite trends: Gemini 1.5 Pro had a slight increase in performance when using plots, while Gemini 1.5 Flash had a slight increase in performance when using text, though the magnitudes of the gains were small.
@@ 4.3 Ablations
We considered a variety of text and plot ablations to confirm if there were any large gains. All ablations were performed on Gemini Pro 1.5. We used the function identification task to test for any performance differences; details and results are reported in Section S.1.3.
@@ 4.4 Cost
Using plots for time-series data can often be more cost-efficient and token-efficient. Token efficiency matters when your context is large and the context-window is limited; for example we needed to downsample our raw signals to fit them into the 128k context window for GPT4o(-mini), particularly with the large few-shot experiments (Section 4.2). For example, when using the Gemini API (Google, 2024), images account for 258 tokens if both dimensions are less than 384 x 384 pixels, after which 4 additional crops are added for a total of 1290 tokens. Text tasks can easily be 10x larger (e.g. 10-shot activity recognition Section 3) requiring more than the entire 128k context available.
Plot experiments also end up being cheaper. As an example, for our most expensive experiment on few-shot activity recognition, we can estimate the input token cost of our 5-shot experiment for both plot and text on GPT4o (OpenAI, 2024). The text version of this task required nearly 128k text tokens (costing \$0.32 per 5-shot question); by contrast, the plot version required 50 images (costing \$0.032), a 10x difference in overall costs for input tokens.
@@ 5 Conclusions and Future Work
Our key finding is that engaging the vision encoder of a multimodal foundation model through the use of plot representations leads to significant performance and efficiency gains on time-series understanding tasks, compared with relying on the text encoder. By processing data visually instead of textually, these models can better capture temporal patterns and relationships. We established our results on synthetic data with well-controlled characteristics and reasoning types, and also showed that this approach holds on real, noisy and complex tasks related to making sense of consumer health signals.
The method presented here is powerful in its simplicity and generalizability – we believe that it is particularly useful when the following conditions are met:
  • You want to use an off-the-shelf multimodal model to interpret your time-series data.
@@
All plots in this work were generated by human-written code in order to avoid any bias. Looking forward, in real applications we anticipate that plotting could be part of a tool-use framework, where the model is prompted to choose how and when to plot the data, after which it uses the plot representation it created.

### v2

@@ 1 Introduction
Multimodal models like GPT4 (Achiam et al., 2023) and Gemini (Gemini Team et al., 2023) are trained to understand visual information natively. However, they are not specifically trained to understand time-series data – in particular, the tokenizers for large language models (LLMs) are not well-suited for representing large sequences of floating point numbers (Spathis & Kawsar, 2024). This mirrors the human approach (Card et al., 1999; Yalçin et al., 2016); we cannot easily make sense of a long array of floating point numbers - instead our first instinct is often to visualize the data through plotting, followed by extracting insights through statistical analysis.
@@
Our work empirically evaluates the relative performance of the native capabilities of existing multimodal foundation models on visual versus textual representations of time-series data. This contribution furthers the understanding of modern foundation models, which is important in real-world contexts as user-facing products continue to develop multimodal sophistication and users interrogate data types that are more complex than can be easily represented with text only. While we do not claim to achieve the same absolute performance as models trained for specific tasks, and likely our approach will not match such models, our results presented here nonetheless show that in contexts when one relies on a foundation model to ingest any general time-series data with a priori unknown characteristics, a visual representation is on balance likely to yield better, and cheaper, results.
@@ 2 Related Work
We use the term “time-series understanding” in this paper to distinguish from time-series forecasting. Time-series forecasting predicts future data points based on points seen so far, whereas we are primarily interested in the setting where we connect the time-series data to a multimodal model for further analysis. In particular, we want to show that multimodal models can reason about overall trends, the relationship between multiple time-series, overall clustering of data, and other time-series understanding tasks. Forecasting has been a very productive area of the field – for a closer look at the literature, the survey paper by Zhang et al. (2024) tracks a wide range of time-series understanding and forecasting work.
Time-series models
@@
There are several existing approaches that train time-series encoders for specific tasks or domains. For example, Chan et al. (2024) trained a domain-specific encoder for multimodal medical time-series analysis. Similarly, Cosentino et al. (2024) trained an encoder for the sleep data in their Digital Well-being task. In this paper, we are not claiming that our approach would outperform a task-specific model on a specific task – we claim that one can achieve much better performance from a foundation model by exploiting its native multimodal capabilities compared to using only text.
Others have also shown that training foundation models with Transformer architectures specifically to work in time-series contexts can lead to good results across tasks including mostly forecasting but also classification, anomaly detection and imputation. These trained models do especially well when the input time-series are carefully pre-processed and tokenized, including patching, scaling and quantization Das et al. (2023); Woo et al. (2024); Goswami et al. (2024); Ansari et al. (2024); Cai et al. (2023). While they did not train a new model, in LLMTime (Gruver et al., 2024) the authors showed that with careful tokenization, text-only LLMs can perform well at forecasting tasks; we perform ablations based on their methods and select the best tokenizations accordingly for our text baselines.
In this work, our goal is to show that simply plotting the data without additional data preprocessing or model training is at the very least an easy first step, and might be a helpful approach when training a task-specific encoder from scratch may not be feasible due to the requirements on having additional paired data, compute and expertise.
Vision models and visual representations
@@
While studying multimodal models’ abilities to reason about visual inputs, Rahmanzadehgervi et al. (2024) found that multimodal models are unable to reason effectively, although some follow-up work by Corin (2024) indicated that prompt engineering can fix losses. In any case, our results do not necessarily contradict this - for many of our tasks humans may be able to get perfect scores, and indeed the multimodal models do not. Regardless, our main claim that plots are better suited than text as input to a multimodal model for time-series understanding holds true.
Perhaps an inversion of our approach, DePlot (Liu et al., 2023) translated visual plots to numeric tables and operated on the tabular data. This approach may be sound for discrete data where the number of points remains small – cases where a human would be expected to understand a table of data well.
Closely related to our work are those methods that use vision-embedding models like Contrastive Language–Image Pre-training (CLIP) by Radford et al. (2021). Wimmer & Rekabsaz (2023) used CLIP to embed plots of financial time-series data, from which features are extracted for use by downstream classifiers. Since we use the vision encoders of the multimodal foundation models directly, there is no need for further feature extraction or downstream classifiers in our approach. IMU2CLIP (Moon et al., 2023) and ImageBind (Girdhar et al., 2023) used video and image data paired with waveforms to learn joint embeddings. Both of these works rely on existing paired waveform and video data to “bind” the modalities together, whereas we can simply plot the waveforms and use the existing multimodal vision encoder to derive our time-series embeddings.
Measuring understanding
Past work has also investigated various approaches to measuring the degree to which models can understand and reason about various modalities of inputs, such as charts (CharXiv (Wang et al., 2024)), tables and figures (SPIQA (Pramanick et al., 2024)) and time-series themselves (TimeSeriesExam (Cai et al., 2024)). These works generally involve generating novel evaluation datasets, and in some cases (e.g. (Wang et al., 2024), (Pramanick et al., 2024)) rely on language models to generate the questions themselves. In our work we deliberately avoid this approach as it can introduce biases during evaluation that are hard to account for (e.g. favoring their own output as in (Panickssery et al., 2024)). In TimeSeriesExam (Cai et al., 2024), the authors also found, as we do, that models perform better on plot-based representations of time-series than the analogues text-based representations, though only demonstrate this on synthetic data as part of a carefully optimised exam generation algorithm.
@@ 3 Methodology
We evaluate our visual prompting method on both synthetic data and real-world use-cases. Synthetic data allows us to control the difficulty of the task by adding noise and altering the number of data points, and to investigate specific kinds of reasoning in isolation. We chose the synthetic tasks to align with the different steps of reasoning we hypothesize are required for the representative real-world use cases we test on. Note that in this context, “reasoning” refers to the high-level steps we believe humans would take to get to the right answer, rather than any formal modelling approach such as chain of thought. These tasks include understanding the local and global longitudinal signatures (trend and magnitude) of a time-series, and potentially comparing it with several other time-series (as in the case of multidimensional sensing). We summarize in Section 4.1 the different tasks in our experiments, along with the type of reasoning we are probing.
The goal of our work is to study specifically the difference in performance achieved by models when ingesting visual versus textual representations rather than the absolute performance on any one task with either modality. As such the appropriate baseline, and the one we use, is the models’ performances on textual representations. We nonetheless include random choice baselines for context to show there is utility in leveraging these models at all, and compare against state-of-the-art task-specific models (for two of the real-world tasks) for context.
@@ 4 Experiments and Results
Type of data Task Reasoning Time-series length
@@
Table 1: Summary of the tasks we study in this paper, including the reasoning each requires and the length of the input time-series (based on the number of points). The “reasoning” column is a high-level summary of the steps needed to answer the tasks’ questions correctly; tasks are detailed in Sections 4.1 and 4.2 and Supplementary Information Section S.1.1.
@@
Few-Shots GPT4o-mini Gemini Flash 1.5 GPT4o Gemini Pro 1.5
@@
Table 2: Summary of the experiments with 38 out of 42 results showing better performance on plots (bold numbers). Cells contain metric medians and IQRs. Stars in synthetic tasks only indicate statistically significant differences between plot and text metrics at 95% confidence corrected for multiple comparisons; we could not perform the same hypothesis testing on the real-world tasks. See Section 3.4 for statistical details and Supplementary Table S1 for relative differences between approaches.
@@ 4.1 Synthetic Data Tasks
Figure 1 summarizes all the zero-shot versions of the synthetic data tasks showing that plot-based methods outperform the text-based methods across GPT and Gemini model families, with few exceptions. Detailed task descriptions are provided in Supplementary Information Section S.1.1.1, and non-aggregated results over dataset parameters such as number of points and noise level are available in Supplementary Information Section S.1.2.
@@
Figure 1: Zero-shot synthetic data results showing plot- and text-based accuracy (MAE for the cluster counting task) distributions for all models, with horizontal lines representing random performance. The results generally show better performance for plots compared to text across models.
@@
This is the simplest task that requires only identifying one overall trend and correctly classifying it into one of five functional tasks (linear, quadratic, cubic, exponential or periodic). We generate a set number of points with a controlled level of noise according to one of the five function classes, and test the model’s ability to label the global trend into the correct class.
@@
This task now requires understanding the trends in two lines and comparing them against each other to correct identify whether the lines are positively or negatively correlated. Here we generate two linear series with controlled number of points and noise and predetermined slopes, measure the sign of the correlation analytically using the Pearson correlation coefficient, and probe the model’s ability to identify the directional correlation.
@@
In this task, the model needs to correctly identify the $N$ number of clusters present in a set of points. To test this, we generate random points on a 2D grid with a set number of clusters and controlled minimum distance between cluster centers and standard deviation of the points about the centres. The model is then instructed to count the number of clusters.
@@
This is a harder task: the model must now identify the correct first derivative (out of four choices) of the function provided in the question. The function and choices are presented as either plots or text. We pass various known functions to the model alongside four synthetic first derivatives, each corresponding to different functional classes, with controlled levels of noise and number of points. The models are then asked to identify which of the multiple choices corresponds to the true derivative.
@@
As a hard variant of derivative identification, we always pass a quadratic function and present four different linear functions as multiple choice answers, with controlled levels of noise and number of points. The model must now identify the correct linear function (out of four choices) that corresponds to the quadratic function’s derivative, so it must pay attention to both the sign and magnitude of the linear slopes.
@@
Figure 2: Quadratic derivative identification results show zero-shot plots outperform text, except for the outlier GPT4o model. When using few-shots, more examples generally improves the gain.
@@ 4.2 Real-world Tasks
Our synthetic tasks built up in complexity from understanding one trend globally (functional form identification) to understanding several trends simultaneously using global and local information (correlation and cluster counting) and complex multi-step reasoning (derivative identifications). We now probe tasks on real-world data that require a mix of these reasoning abilities, including simultaneous understanding of multiple sensor signals to uncover either local or global trends in the first two tasks (fall detection and activity recognition), and extracting two trends of different timescales in the last task (readiness).
Figure 3: Results of fall detection task show consistently better plot performances across models and number of few-shots, with plot performance generally increasing with number of shots. The top plot models have 10-shot (sensitivity, specificity) as follows: Gemini Pro 1.5 - (0.84, 0.95) and GPT4o - (0.92, 0.81), compared to the state-of-the-art task-specific support-vector machine model reported by Aziz et al. (2017) which achieves (0.96, 0.96) (see Supplementary Table S2 for more details).
@@
An IMU segment is a 6D-vector composed of 3-axes accelerometer signals and 3-axes gyroscope signals. The first real-world task we evaluate (results in Figure 3) is to classify whether a 15-second IMU segment recorded at 128hz contains a fall, a “near” fall or showed “active daily living” (ADL). The dataset used in the open-source IMU Fall Detection dataset (IMUFD, Aziz et al. (2017)).
@@
Figure 4: Results of activity recognition task for all models across few-shot numbers (where context length allowed), showing overall improved performance for plots. The performance of the state-of-the-art deep-learning model reported by Kumar & Selvam (2022) is included for reference.
@@
A further real-world IMU task we evaluate is to classify whether a 15-second IMU segment is one of five activity classes: “sit”, “stand”, “stairs”, “walk” or “bike” (results in Figure 4). The dataset used in the open-source Heterogeneity Human Activity Recognition dataset (HHAR, Stisen et al. (2015)). As with Fall Detection we test performance with 1, 3, 5 and 10 few-shot examples, with 383 samples per model and number of few-shots. For this task the text representation of the 10 few-shot examples exceeded the GPT4o and GPT4o-mini 128k context windows, so these results are only shown for the Gemini models.
@@
Figure 5: Results of readiness task for Gemini models only (as the dataset cannot be sent to other models), demonstrating approximate parity between the text and plot approaches.
@@
We use training load data from 350 fitness case studies (Cosentino et al., 2024) and present it as tables or TRIMP bar plots (results shown in Figure 5). Each case study contains data from 30 consecutive days. We use a simplified version of the textual prompt and visualization from Cosentino et al. (2024) and do not split TRIMP in different heart rate zones. We tested Gemini 1.5 Pro and Gemini 1.5 Flash for both plot and text approaches as zero-shot tasks. Since this task involved analyzing just 30 data points we did not expect the plot prompt to excel here. Interestingly, models of different sizes showed opposite trends: Gemini 1.5 Pro had a slight increase in performance when using plots, while Gemini 1.5 Flash had a slight increase in performance when using text, though the magnitudes of the gains were small.
@@ 4.3 Ablations
We considered a variety of text and plot ablations to confirm if there were any large gains. All ablations were performed on Gemini Pro 1.5. We used the function identification task to test for any performance differences; details, results and visualizations are reported in Section S.1.3.
@@ 4.4 Cost
Using plots for time-series data can often be more cost-efficient and token-efficient. Token efficiency matters when your context is large and the context-window is limited; for example we needed to downsample our raw signals to fit them into the 128k context window for GPT4o(-mini), particularly with the large few-shot experiments (Section 3). For example, when using the Gemini API (Google, 2024), images account for 258 tokens if both dimensions are less than 384 x 384 pixels, after which 4 additional crops are added for a total of 1290 tokens. Text tasks can easily be 10x larger (e.g. 10-shot activity recognition Section 4) requiring more than the entire 128k context available. Depending on the task it may be possible to reduce the number of text tokens by further sub-sampling of the data, but this may result in reduced task performance. The optimal sampling rate may also be task- and dataset-specific.
Plot experiments also end up being cheaper. As an example, for our most expensive experiment on few-shot activity recognition, we can estimate the input token cost of our 5-shot experiment for both plot and text on GPT4o (OpenAI, 2024). The text version of this task required nearly 128k text tokens (costing \$0.32 per 5-shot question); by contrast, the plot version required 50 images (costing \$0.032), a 10x difference in overall costs for input tokens. In addition to being cheaper, the plotting approach also scales better for longer time-series, as the number of tokens required for the textual approach grows with the number of data points while a plot of the same size will generate the same number of tokens independent of the number of data points being plotted.
@@ 5 Conclusions and Future Work
The key finding from our rigorous empirical evaluation is that engaging the vision encoder of a multimodal foundation model through the use of plot representations leads to significant performance and efficiency gains on time-series understanding tasks, compared with relying on the text encoder. By processing data visually instead of textually, these models can better capture temporal patterns and relationships. We established our results on synthetic data with well-controlled characteristics and reasoning types, and also showed that this approach holds on real, noisy and complex tasks related to making sense of consumer health signals. This is analogous to the gains that humans benefit from when visualising data (Card et al., 1999; Yalçin et al., 2016), though we do not claim that the mechanisms by which visual data are processed by the models we study here are the same as those with which humans process visual stimuli.
The method presented here is powerful in its simplicity and generalizability and relies on the native capabilities of multimodal models requiring no additional training – we believe that it is particularly useful when the following conditions are met:
  • You want to use an off-the-shelf multimodal model to interpret your time-series data, as might be the case in user-facing applications that rely on general models to understand a broad range of potential user inputs including natural language.
@@
In this work, all plots were generated by human-written code in order to avoid any bias. As such we do not rigorously study what the optimal plotted form of a certain time-series might be for visual understanding; this is likely to be a function of the exact downstream task or user request, but could in theory be automated and forms the basis for future work. Looking forward, in real applications we anticipate that plotting could be part of a tool-use framework, where the model is prompted to choose how and when to plot the data, after which it uses the plot representation it created.
Lastly, further work remains in the explainability context – while we demonstrate empirically that visual representations generally outperform textual representations of time-series data, we have not probed why mechanistically this is so.